# FDGaussian: Fast Gaussian Splatting from Single Image via Geometric-aware Diffusion Model

### v1

@@ Abstract
Reconstructing detailed 3D objects from single-view images remains a challenging task due to the limited information available. In this paper, we introduce FDGaussian, a novel two-stage framework for single-image 3D reconstruction. Recent methods typically utilize pre-trained 2D diffusion models to generate plausible novel views from the input image, yet they encounter issues with either multi-view inconsistency or lack of geometric fidelity. To overcome these challenges, we propose an orthogonal plane decomposition mechanism to extract 3D geometric features from the 2D input, enabling the generation of consistent multi-view images. Moreover, we further accelerate the state-of-the-art Gaussian Splatting incorporating epipolar attention to fuse images from different viewpoints. We demonstrate that FDGaussian generates images with high consistency across different views and reconstructs high-quality 3D objects, both qualitatively and quantitatively. More examples can be found at our website https://qjfeng.net/FDGaussian/.
Keywords:
3D Reconstruction Gaussian Splatting Diffusion Model
Figure 1: Given a single reference image as input, our method can obtain geometric details while maintaining multi-view consistency.
@@ 1 Introduction
Single-view 3D reconstruction aims to recover 3D geometry and appearance of an object from a single RGB image. This task holds immense importance as it allows machines to understand and interact with the real 3D world, enabling various applications in virtual reality (VR), augmented reality (AR)[22, 24] and robotics[53].
A major challenge in 3D reconstruction involves securing a representation that is both high-quality and efficient. Explicit representations like point clouds[38, 39, 1, 13], voxels[33, 10, 65, 61], and meshes[15, 52] are commonly used due to their intuitive and deformation friendly property, but struggle to represent realistic appearance. In recent years, implicit representations (e.g., Neural Radiance Field (NeRF)[36, 27, 69]) have witnessed great success since the continuous nature of these methods helps optimization. However, the stochastic sampling required for rendering is time-consuming and may result in noise.
@@
To leverage the strength of both implicit and explicit radiance fields while overcoming their drawbacks, 3D Gaussian Splatting[21] is proposed. It combines the benefits of neural network-based optimization and explicit, structured data storage, allowing for high-quality rendering with competitive training and inference time. Current methods[49, 70, 50] often feed a single image for Gaussian Splatting, ignoring the spatial correspondence of multiple views. Additionally, we observe that the original implementation of Gaussian Splatting[21] neglects the distance between 3D Gaussians, causing many unnecessary split and clone operations.
Several works[51, 32] have shown that fine-turning a 2D generator to understand the correlation between different views of the object significantly facilitates 3D reconstruction. Yet these methods either suffer from multi-view inconsistency or struggle to handle objects with complicated geometric structures.
@@
Considering these, we propose FDGaussian, a novel two-stage framework for single-image 3D reconstruction composed of a geometric-aware multi-view generation stage and the following accelerated 3D Gaussian reconstruction stage. The generation stage aims to synthesize 3D-aware and multi-view consistent high-fidelity images. To achieve the goal, 3D features are extracted as geometric condition by decoupling the orthogonal planes while semantic condition is obtained with the CLIP[40] encoder. Together with the input image, both conditions are fed into the diffusion model[44]. At the reconstruction stage, we introduce epipolar attention to fuse the generated consistent views, which fully exploits the underlying geometric correlation between views, allowing for competitive visual quality enhancement. Moreover, to further accelerate the optimization process, we propose a novel metric termed Gaussian Divergent Significance (GDS) to avoid unnecessary operations.
@@
Extensive experiments and ablations on Objaverse[11] and GSO[12] dataset demonstrate that our method is able to generate high-quality 3D objects with multi-view consistency and detailed geometric. Furthermore, we show that FDGaussian can seamlessly integrate with text-to-image models in downstream text-to-3D applications. Our main contributions can be summarized as following:
@@
  • In order to take full advantage of the consistent multi-view images, we introduce epipolar attention into the rendering process, allowing for efficient and effective communication between images.
  • We derive a novel metric named Gaussian Divergent Significance (GDS) to prune unnecessary split and clone operations during optimization, achieving significant time reduction.
@@ 2 Related Work
2.1 Representations for 3D Reconstruction
2.1.1 Explicit representations
Explicit representations have been predominant in industries and academic research for a long time. Classic representations, including point clouds[38, 13, 35], voxels[33, 10, 59, 61, 65, 60], meshes[58, 66, 52, 55], have been revisited for 3D reconstruction. While these explicit representations offer detailed descriptions of geometry and appearance, they lack the flexibility of underlying topology and often struggle to capture realistic appearances effectively.
2.1.2 Implicit representations
Different from explicit representations, implicit representations, including signed distance fields (SDF)[35, 6, 8] and unsigned distance fields (UDF)[9, 31, 16], offer the advantage of accurately modeling arbitrary geometry and topology. Thanks to the continuous nature of implicit representations, they can leverage deep neural networks to support data-driven geometry learning. In recent years, Neural Radiance Field (NeRF)[36] has demonstrated encouraging progress, allowing for 3D optimization with only 2D supervision via volumetric rendering. Nevertheless, implicit approaches suffer from extensive sampling to fit the implicit functions of 3D scenes. This leads to significant computational costs, particularly in high-resolution or interactive rendering scenarios, even with accelerated NeRF versions[5, 2, 14, 45]. It is difficult to achieve real-time rendering and high-quality view synthesis at the same time.
2.1.3 Gaussian splatting
In order to tackle the aforementioned obstacles, Gaussian Splatting[21] has emerged as an alternative representation and has shown remarkable advancements in terms of both quality and speed, offering a promising avenue. Some methods[50, 70] leverage Gaussians to generate coarse 3D representation and then refine it with reference information, which often requires extra depth input or mesh extraction process. Other methods[49] mitigate this problem by directly predicting 3D representation from the monocular input image, yet suffering from artifacts in unseen regions. In contrast to these works, our method combines the creativity of diffusion models and the efficiency of Gaussian Splatting and can render high-quality Gaussian representation in a relatively short time without explicit depth or normal hint.
2.2 3D Reconstruction Guided by 2D Diffusion Models
The recent success of denoising diffusion probabilistic models (DDPM)[46, 18, 47, 48, 64, 62, 63] has caught a surge of interest. In visual content creation, language-guided image diffusion models such as DALL-E2[43], Imagen[42], and Stable Diffusion[44] have shown huge potential in generating photorealistic images with strong semantic correlation to the given text-prompt inputs. They possess powerful priors about our 3D world and have inspired a growing body of research to employ 2D prior models for assisting 3D generative tasks. Many methods follow the paradigm of per-shape optimization[25, 3, 19, 37, 7, 34, 23]. They typically optimize a 3D representation and utilize 2D diffusion models for gradient guidance. While they have yielded impressive results, these methods tend to suffer from prolonged optimization times, the "multi-face" problem, over-saturated colors, and a lack of diversity in results.
A new wave of studies, highlighted by works like Zero-1-to-3[30], has showcased the promise of using pre-trained 2D diffusion models for synthesizing novel views from singular images or text, opening new doors for 3D generation. Nevertheless, the multi-view images produced by previous methods[30, 29, 68] lack consistency and geometric details. Our research, along with several concurrent studies[28, 26, 56, 4, 32, 57], is dedicated to improving multi-view consistency while capturing the intricate structure of the reference image.
3 Method
@@
3.1 Geometry-aware Multi-view Image Generation
Finetuning pre-trained diffusion models [44] to synthesize novel images under a given camera transformation has demonstrated promising results[30, 32, 57, 68]. One stream of methods[57, 68] address the multi-view inconsistency problem by conditioning on previously generated images, which tends to be susceptible to cumulative errors and reduced processing speeds. Another stream of methods[30, 28] solely use the reference image and semantic guidance to generate novel views, but suffer from collapsed geometry and limited fidelity.
We argue that the pivot lies in fully utilizing the geometric information provided by the reference image. However, directly extracting 3D information from a single 2D image is not feasible. Thus, it is imperative to effectively disentangle 3D features from the image plane (i.e. $xy$-plane) by decoupling orthogonal planes. We first employ a vision transformer to encode the input image and capture overall correlations in the image, generating high-dimensional latent $\boldsymbol{h}$. Then we leverage two decoders, an image-plane decoder and an orthogonal-plane decoder, to generate geometric-aware features from the latent. The image-plane decoder reverses the encoding operation, leveraging a self-attention mechanism on the encoder output and converting it into $F_{xy}$. In order to generate orthogonal-plane features while maintaining structural alignment with the image plane, a cross-attention mechanism is employed to decode $yz$ and $xz$ plane features $F_{yz}$ and $F_{xz}$. To facilitate the decoding process across different planes, we introduce a learnable embedding $\boldsymbol{u}$ that supplies additional information for decoupling new planes. The learnable embedding $\boldsymbol{u}$ is first processed through self-attention encoding and then used as a query in a cross-attention mechanism with the encoded image latent $\boldsymbol{h}$. The image features are converted into keys and values for the cross-attention mechanism as following:
$\texttt{CrossAttn}(\boldsymbol{u},\boldsymbol{h})=\texttt{SoftMax}\bigg{(}\dfrac{(W^{Q}\texttt{SelfAttn}(\boldsymbol{u}))(W^{K}\boldsymbol{h})^{T}}{\sqrt{d}}\bigg{)}(W^{V}\boldsymbol{h}),$ (1)
where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable parameters and $d$ is the scaling coefficient. Finally, the features are combined as geometric condition:
$F=F_{xy}\textcircled{c}(F_{yz}+F_{xz}),$ (2)
where c⃝ and $+$ are concatenation and summation operations, respectively.
Backbone design. Similar to previous works[44, 18], we use a latent diffusion architecture with an encoder $\mathcal{E}$, a denoiser UNet $\epsilon_{\theta}$, and a decoder $\mathcal{D}$. The network is initialized from the pre-trained weights of Zero-1-to-3[30] given its massive scale of training data. Following [30] and [32], the input view is channel-concatenated with the noisy target view as the input to UNet. We employ the CLIP image encoder[40] for encoding $\mathcal{I}_{ref}$, while the CLIP text encoder[40] is utilized to encode $\Delta\pi$. The concatenation of their embeddings, denoted as $c(\mathcal{I}_{ref},\Delta\pi)$, forms the semantic condition in the framework. We can learn the network by optimizing the following objective:
$\mathop{\min}_{\theta}\mathbb{E}_{z\sim\mathcal{E}(\mathcal{I}),t,\epsilon\sim\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(z_{t},t,c(\mathcal{I}_{ref},\Delta\pi))\|_{2}^{2}$ (3)
3.2 Preliminary of Gaussian Splatting
@@
3D Gaussian Splatting is a learning-based rasterization technique for 3D scene reconstruction and novel view synthesis[21]. Each Gaussian element is defined with a position (mean) $\boldsymbol{\mu}$, a full 3D covariance matrix $\boldsymbol{\Sigma}$, color $c$, and opacity $\sigma$. The Gaussian function $G(x)$ can be formulated as:
$G(x)=exp(-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})).$ (4)
@@
To ensure the positive semi-definiteness of $\boldsymbol{\Sigma}$, the covariance matrix $\boldsymbol{\Sigma}$ can be factorized into a scaling matrix $S$ represented by a 3D-vector $s\in\mathbb{R}^{3}$ and a rotation matrix $R$ expressed as a quaternion $q\in\mathbb{R}^{4}$ for the differentiable optimization: $\boldsymbol{\Sigma}=RSS^{T}R^{T}$.
@@
The rendering technique of splatting, as initially introduced in [21], is to project the Gaussians onto the camera image planes, which are employed to generate novel view images. Given a viewing transformation $W$, the covariance matrix $\boldsymbol{\Sigma}{{}^{\prime}}$ in camera coordinates is given as: $\boldsymbol{\Sigma}{{}^{\prime}}=JW\boldsymbol{\Sigma}W^{T}J^{T}$, where $J$ is the Jacobian matrix of the affine approximation of the projective transformation. After mapping 3D Gaussians to a 2D image space, we count 2D Gaussians that overlap with each pixel and calculate their color $c_{i}$ and opacity $\sigma_{i}$ contribution. Specifically, the color of each Gaussian is assigned to every pixel based on the Gaussian representation described in Eq. 4. And the opacity controls the influence of each Gaussian. The per-pixel color $\hat{C}$ can be obtained by blending N ordered Gaussians: $\hat{C}=\sum_{i\in N}c_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{i})$.
Figure 3: Qualitative results of novel view synthesis and 3D reconstruction. The nearby views maintain high fidelity to the reference image, while the distant views generate a variety of expressive yet realistic images.
3.3 Accelerating the Optimization
However, the split and clone operations proposed by the original work[21] overlook the distance between 3D Gaussians, during the optimization process which significantly slows down the process. We observe that if two Gaussians are close to each other, even if the positional gradients are larger than a threshold, they should not be split or cloned since these Gaussians are updating their positions. Empirically splitting or cloning these Gaussians has negligible influence on the rendering quality as they are too close to each other. For this reason, we propose Gaussian Divergent Significance (GDS) as a measure of the distance of 3D Gaussians to avoid unnecessary splitting or cloning:
$\Upsilon_{GDS}(G(\boldsymbol{x}_{1}),G(\boldsymbol{x}_{2}))=\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}+tr(\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}-2(\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\Sigma}_{2}\boldsymbol{\Sigma}_{1}^{-1})^{1/2}),$ (5)
where $\boldsymbol{\mu}_{1}$, $\boldsymbol{\Sigma}_{1}$, $\boldsymbol{\mu}_{2}$, $\boldsymbol{\Sigma}_{2}$ are the position and covariance matrix of two 3D Gaussians $G(\boldsymbol{x}_{1})$ and $G(\boldsymbol{x}_{2})$.
In this way, we only perform the split and clone operations on the 3D Gaussians with large positional gradients and GDS. To avoid the time-consuming process of calculating GDS for every pair of 3D Gaussians, we further propose two strategies. Firstly, for each 3D Gaussian, we locate its closest 3D Gaussian by leveraging the k-nearest neighbor (k-NN) algorithm and calculate their GDS for each pair. As a result, the time complexity is reduced from $O(N^{2})$ to $O(N)$. Additionally, as mentioned in Sec. 3.2, the covariance matrix can be factorized into a scaling matrix $S$ and a rotation matrix $R$: $\boldsymbol{\Sigma}=RSS^{T}R^{T}$. We take advantage of the diagonal and orthogonal properties of rotation and scaling matrices to simplify the calculation of Eq. 5. Details of GDS will be discussed in the Supplementary Materials.
3.4 Epipolar Attention for Multi-view Rendering
Previous methods[50, 70] usually use a single input image for coarse Gaussian Splatting, which requires further refinement or repainting in unseen regions. The intuitive idea is to exploit the generated consistent multi-view images to reconstruct high-quality 3D objects. However, relying solely on cross-attention to communicate between images of multiple viewpoints is insufficient. Therefore, given a sequence of generated views, we propose epipolar attention to allow association between the features of different views. The epipolar line for a given feature point in one view is the line on which the corresponding feature point in the other view must lie, based on the known geometric relationship between two views. It acts as a constraint to reduce the number of potential pixels in one view that can attend to another view. We present the illustration of epipolar line and epipolar attention in Fig. 4. By enforcing this constraint, we can limit the search space for corresponding features in different views, making the association process more efficient and accurate.

### v2

# GeoGS3D: Single-view 3D Reconstruction via Geometric-aware Diffusion Model and Gaussian Splatting

@@ Abstract
We introduce GeoGS3D, a novel two-stage framework for reconstructing detailed 3D objects from single-view images. Inspired by the success of pre-trained 2D diffusion models, our method incorporates an orthogonal plane decomposition mechanism to extract 3D geometric features from the 2D input, facilitating the generation of multi-view consistent images. During the following Gaussian Splatting, these images are fused with epipolar attention, fully utilizing the geometric correlations across views. Moreover, we propose a novel metric, Gaussian Divergence Significance (GDS), to prune unnecessary operations during optimization, significantly accelerating the reconstruction process. Extensive experiments demonstrate that GeoGS3D generates images with high consistency across views and reconstructs high-quality 3D objects, both qualitatively and quantitatively. Further examples can be found at the website.
Figure 1: GeoGS3D for single image to 3D generation: GeoGS3D can reconstruct 3D content with detailed geometry and accurate appearance from a single image.
@@ 1 Introduction
Single-view 3D reconstruction aims to recover 3D geometry and appearance of an object from a single RGB image. This task holds immense importance as it allows machines to understand and interact with the real 3D world, enabling various applications in virtual reality (VR), augmented reality (AR) (Kopf et al., 2020; Li et al., 2023) and robotics (Wang et al., 2022). Despite broad applications of the task, it remains highly challenging due to the inherent complexity of inferring 3D structures from a single 2D image.
Poineering works in multi-view diffusion models have made strides in improving 3D reconstruction by finetuning pre-trained image or video diffusion models on 3D datasets to enable multi-view synthesis (Liu et al., 2023c; Shi et al., 2023; Kwak et al., 2024; Huang et al., 2024; Voleti et al., 2024; Long et al., 2024). However, maintaining multi-view consistency and handling objects with complex geometries remain difficult. Another line of research (Hong et al., 2023; Tang et al., 2024a; Xu et al., 2024c; b) proposes generalizable reconstruction models, generating 3D representation from one or few views in a feed-forward process, but these models usually demand substantial computational resources. For instance, LGM (Tang et al., 2024a) requires training on 32 NVIDIA A100 (80G) for 4 days.
@@
Meanwhile, 3D Gaussian Splatting (Kerbl et al., 2023) has emerged as a promising technique, offering high-quality rendering with competitive training and inference time as it combines the benefits of neural network-based optimization and explicit, structured data storage. Hence, it is desirable to explore Gaussian Splatting for efficient 3D reconstruction. However, current methods (Szymanowicz et al., 2023a; Tang et al., 2023a) using 3D Gaussian Splatting often feed a single image into the model, ignoring the spatial correspondence of multiple views. Additionally, we observe that the original implementation of Gaussian Splatting neglects the distance between 3D Gaussians, causing many unnecessary split and clone operations.
@@
To address these limitations, we propose GeoGS3D, a novel two-stage framework for single-image 3D reconstruction. Our approach consists of a geometric-aware multi-view generation stage, followed by an accelerated 3D reconstruction stage. In the generation stage, we aim to synthesize 3D-aware images that maintain multi-view consistency. To achieve the goal, 3D features are extracted as geometric conditions by decoupling the orthogonal planes, while the semantic condition is obtained with the CLIP (Radford et al., 2021) encoder. These conditions, together with the input image are then fed into the diffusion model (Rombach et al., 2022). In the reconstruction stage, we introduce epipolar attention to fuse the generated views, fully leveraging the geometric correlations between them. Furthermore, we accelerate the optimization process by introducing a novel metric, Gaussian Divergent Significance (GDS), to avoid unnecessary operations during 3D Gaussian Splatting.
@@
Extensive experiments and ablations on Objaverse (Deitke et al., 2023) and Google Scanned Object (Downs et al., 2022) datasets demonstrate that our method is able to produce high-quality 3D objects with strong multi-view consistency and detailed geometry. Our key contributions can be summarized as following:
@@
  • In order to take full advantage of the consistent multi-view images, we introduce epipolar attention into the optimization process, allowing for efficient and effective communication between images.
  • To accelerate the optimization of Gaussian splatting, we derive a novel metric named Gaussian Divergent Significance (GDS) to prune unnecessary split and clone operations during optimization.
@@ 2 Related Work
2.1 Multi-view Diffusion Models
Recent 2D diffusion models (Song et al., 2020a; Ho et al., 2020; Song & Ermon, 2019; Song et al., 2020b; Xing et al., 2023c; a; b) make impressive advances in generating images from various conditions. As an intermediate 3D representation, the generation of multi-view images using diffusion models has been explored. The advantage of multi-view images is that they are batched 2D projections and can be directly processed by existing image diffusion models with minor changes. Zero-123 (Liu et al., 2023c) injects camera view as an extra condition to the diffusion model for generating images from different perspectives. Additionally,  Shi et al. (2023) proposes replacing self-attention with multi-view attention in the UNet to produce multi-view consistent images. Other works (Chan et al., 2023; Liu et al., 2023a; Szymanowicz et al., 2023b; Yang et al., 2024; Tseng et al., 2023; Sargent et al., 2023) follow a similar approach to make diffusion model 3D-aware and enhance generation consistency. However, most of these works are not designed for reconstruction tasks and still rely on SDS loss to derive 3D content. Recent efforts such as SyncDreamer (Liu et al., 2023d) and Wonder3D (Long et al., 2024) generate multi-view consistent 2D representations, subsequently applying reconstruction methods to obtain 3D content. Despite these advancements, the resulting 3D outputs often lack comprehensive geometric information. In contrast, our method generates more consistent views with detailed geometric information, facilitating high-quality 3D content reconstruction in terms of texture and geometry.
2.2 3D Reconstruction Pipelines
Researchers have explored directly training diffusion models on 3D representations (Jun & Nichol, 2023; Müller et al., 2023; Shue et al., 2023). However, they require exhaustive 3D data and computation resources and are also limited to category-level shape generation with simple textures. Other works (Poole et al., 2022; Melas-Kyriazi et al., 2023; Kwak et al., 2024; Sun et al., 2023; Tang et al., 2023a; b; Wang et al., 2023; Zhang et al., 2023) proposed to lift 2D diffusion models for 3D generation via Score Distillation Sampling (SDS), optimizing various 3D representations including NeRF, mesh, SDF and Gaussian Splatting. However, these approaches often face challenges such as prolonged optimization times, the Janus problem, and over-saturated colors. More recent works (Xu et al., 2023; Szymanowicz et al., 2023b; Siddiqui et al., 2024; Charatan et al., 2024; Xu et al., 2024a; Tochilkin et al., 2024; Li et al., 2024) train large reconstruction models to directly map sparse views to 3D representations. Notably, LRM (Hong et al., 2023), Instant3D (Li et al., 2023), and GRM (Xu et al., 2024c) utilize large transformers to predict triplanes from single or sparse views. TriplaneGaussian (Zou et al., 2024) and LGM (Tang et al., 2024a) instead map sparse views into more memory-efficient 3D Gaussian Splatting, allowing for much higher resolution in supervision. Our work aligns with Liu et al. (2023b; a); Tang et al. (2024b); Wen et al. (2024), where we first multi-view diffusion model to generate consistent images, and then utilize reconstruction methods to obtain the corresponding 3D model.
3 Methodology
@@
The overview of GeoGS3D is shown in Figure 2. Starting from an input image, our geometric-aware diffusion model first generates multi-view images sequentially in the generation stage (refer to Section 3.2). Subsequently, in the reconstruction stage, epipolar attention is incorporated into the Gaussian Splatting process to reconstruct the high-quality 3D objects (refer to Section 3.3). Additionally, we introduce a metric to accelerate the adaptive density control during Gaussian Splatting (refer to Section 3.4).
3.1 Preliminaries
Notations
Given an input image $\mathcal{I}_{0}$, our method aims to generate multiview-consistent images $\{\mathcal{I}_{i}\}_{i=1}^{N}$ and then reconstruct high-quality 3D Gaussians $\mathcal{G}$ with them. The multi-view images are obtained through a diffusion model conditioned on the input image $\mathcal{I}_{0}$ and a set of corresponding relative camera pose change $\{\Delta\pi_{i}\}_{i=1}^{N}$ during the generation stage. In the following reconstruction stage, the 3D Gaussians $\mathcal{G}$ are then optimized with these multi-view images through an accelerated Gaussian Splatting.
3D Gaussian Splatting
@@ 3.1 Preliminaries
(Kerbl et al., 2023) is a learning-based rasterization technique for 3D scene reconstruction and novel view synthesis. Each Gaussian element is defined with a position (mean) $\bm{\mu}$, a full 3D covariance matrix $\bm{\Sigma}$, color $c$, and opacity $\sigma$. The Gaussian function $G(x)$ can be formulated as:
$G(x)=exp(-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})).$ (1)
@@
To ensure the positive semi-definiteness of $\bm{\Sigma}$, the covariance matrix $\bm{\Sigma}$ can be factorized into a scaling matrix $S$ represented by a 3D-vector $s\in\mathbb{R}^{3}$ and a rotation matrix $R$ expressed as a quaternion $q\in\mathbb{R}^{4}$ for the differentiable optimization: $\bm{\Sigma}=RSS^{T}R^{T}$.
@@
The rendering technique of splatting, as initially introduced in Kerbl et al. (2023), is to project the Gaussians onto the camera image planes, which are employed to generate novel view images. Given a viewing transformation $W$, the covariance matrix $\bm{\Sigma}{{}^{\prime}}$ in camera coordinates is given as: $\bm{\Sigma}{{}^{\prime}}=JW\bm{\Sigma}W^{T}J^{T}$, where $J$ is the Jacobian matrix of the affine approximation of the projective transformation. After mapping 3D Gaussians to a 2D image space, we count 2D Gaussians that overlap with each pixel and calculate their color $c_{i}$ and opacity $\sigma_{i}$ contribution. Specifically, the color of each Gaussian is assigned to every pixel based on the Gaussian representation described in Equation 1. And the opacity controls the influence of each Gaussian. The per-pixel color $\hat{C}$ can be obtained by blending N ordered Gaussians: $\hat{C}=\sum_{i\in N}c_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{i})$.
3.2 Geometry-aware Multi-view Generation
Considering the success of pre-trained diffusion models (Rombach et al., 2022), it is intuitive to finetune them for novel view synthesis under a given camera transformation. However, maintaining multi-view consistency across views remains a substantial challenge. One stream of methods (Woo et al., 2023; Ye et al., 2023) addresses the problem by conditioning on previously generated images, which tends to be susceptible to cumulative errors and reduced processing speeds. Another stream of methods (Liu et al., 2023c; a) solely use the reference image and semantic guidance to generate novel views, but suffer from collapsed geometry and limited fidelity.
We argue that the key lies in fully utilizing the geometric information provided by the input image. However, directly extracting 3D information from a single 2D image is not feasible. Thus, it is imperative to effectively disentangle 3D features from the image plane (i.e. $xy$-plane) by decoupling orthogonal planes. We first employ a vision transformer to encode the input image and capture overall correlations in the image, generating high-dimensional latent $\bm{h}$. Then we leverage two decoders, an image-plane decoder and an orthogonal-plane decoder, to generate geometric-aware features from the latent. The image-plane decoder reverses the encoding operation, leveraging a self-attention mechanism on the encoder output and converting it into $F_{xy}$. In order to generate orthogonal-plane features while maintaining structural alignment with the image plane, a cross-attention mechanism is employed to decode $yz$ and $xz$ plane features $F_{yz}$ and $F_{xz}$. To facilitate the decoding process across different planes, we introduce a learnable embedding $\bm{u}$ that supplies additional information for decoupling new planes. The learnable embedding $\bm{u}$ is first processed through self-attention encoding and then used as a query in a cross-attention mechanism with the encoded image latent $\bm{h}$. The image features are converted into keys and values for the cross-attention mechanism as following:
$\texttt{CrossAttn}(\bm{u},\bm{h})=\texttt{SoftMax}\bigg{(}\dfrac{(W^{Q}\texttt{SelfAttn}(\bm{u}))(W^{K}\bm{h})^{T}}{\sqrt{d}}\bigg{)}(W^{V}\bm{h}),$ (2)
where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable parameters and $d$ is the scaling coefficient. Finally, the features are combined as geometric conditions:
[…]
where ⓒ and $+$ are concatenation and summation operations, respectively.
Training Objective
Similar to previous works (Rombach et al., 2022; Ho et al., 2020), we use a latent diffusion architecture with an encoder $\mathcal{E}$, a denoising UNet $\epsilon_{\theta}$, and a decoder $\mathcal{D}$. Following Liu et al. (2023c) and Liu et al. (2023d), the input view is channel-concatenated with the noisy target view as the input to UNet. We employ the CLIP image encoder (Radford et al., 2021) for encoding $\mathcal{I}_{0}$, while the CLIP text encoder (Radford et al., 2021) is utilized for encoding the relative camera pose $\Delta\pi$. The concatenation of their embeddings, denoted as $c(\mathcal{I}_{0},\Delta\pi_{i})$, forms the semantic condition in the framework. We can learn the network by optimizing the following objective:
$\mathcal{L}_{mv}=\mathbb{E}_{z\sim\mathcal{E}(\mathcal{I}),t,\epsilon\sim\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(z_{t},t,c(\mathcal{I}_{0},\Delta\pi_{i}))\|_{2}^{2}$ (4)